OpenStudy (anonymous):

consider the function f(x)=2x^2-8x+5. a)express f(x) in the form of a(x-p)^2+q,where a,p,q are Z? b)find the minimum value of f(x)?

OpenStudy (sriram):

i replied for that jst about an hr ago

OpenStudy (anonymous):

2(x-2)^2-3, min=-3

OpenStudy (jamesj):

First pull out a factor of 2 \[ f(x) = 2(x^2 - 4x) + 5 \] Now you need to 'complete the square'. I.e., what is p such \[ x^2 - 4x = (x-p)^2 + something \]

OpenStudy (anonymous):

x^2-4x+4